OpenStudy (anonymous):

A rocket sled 3 miles in 6 seconds. What is the average speed in miles her hour?

OpenStudy (anonymous):

(3 miles/6seconds)*(3600seconds/1Hour) = 1800miles/hour

OpenStudy (anonymous):

3/6=1/2 seconds. now you have to convert to hours